OpenStudy (anonymous):

Which system has infinite solutions? A) 4x + 6y = -12 2x + 3y = 9 Which system has (-2, -2) as the solution? B) x - y = -2 3x - y = 0 Which system has (1, 3) as the solution? C) x - 4y = -8 2x - 3y = -16 Which system has no solutions? D) x + y = -4 -3x + 2y = 2 Which system has (-8, 0) as the solution? E) x + y = -1 4x + 4y = -4

OpenStudy (anonymous):

2x+3y=9; because you multiply infinitely many ways, x-y=-2; plug it in

OpenStudy (debbieg):

A system will have infinitely many solns if the two equation are REALLY the same equation - that is, the same line. One equation is just a multiple of the other. This is easy to see if you put both equations into y=mx+b form, because the equations will be identical.

OpenStudy (debbieg):

A system has NO solutions if the two lines given by the equations are parallel - because then, then will never intersect, so there is NO point (x,y) that is on both lines. This is easy to see if you put both equations into y=mx+b form, because the SLOPES will be the same, but with different y-intercepts. OR, if you just go about solving by elimination, and you get something where the variables are gone and the equation is clearly FALSE, e.g., 0=5 (or something like that).

OpenStudy (debbieg):

I'm not sure what you're talking about - I don't see that equation in any of these systems?? Your answer is one of the SYSTEMS of 2 equations - A, B, C or D. Not a single equation. I'm not sure where you are getting the equation you state above, because it isn't in any of these systems.

OpenStudy (anonymous):

meant 4x + 6y = -12 2x + 3y = 9

OpenStudy (debbieg):

OK; yes. That's no solution. Like I said: if you use elimination 4x + 6y = -12 (-2)(2x + 3y) = 9(-2) 4x + 6y = -12 -4x - 6y = -18 0=-30 That's false, so it's an inconsistent system - no solution.

OpenStudy (debbieg):

Also, both lines have slope -2/3, but different y-intercepts. So they are parallel.